Find the equations of the spheres which pass through the circle x²+y²+z²=0, x+2y+3z=0 and touch the plane 4x+3y=15
1
Expert's answer
2020-09-15T11:01:52-0400

The equation of a sphere is:

"x²+y²+z²-a-2\lambda(x+2y+3z)=0"

"(x^2-2\lambda x+\lambda^2)-\lambda^2+(y^2-4\lambda y+4\lambda^2)-4\lambda^2+(z^2-6\lambda z+9\lambda^2)-9\lambda^2=a"

"(x-\lambda)^2+(y-2\lambda)^2+(z-3\lambda)^2=a+14\lambda^2"

The radius of the sphere:

"R=\sqrt{a+14\lambda^2}"

The center is:

"(\lambda,2\lambda,3\lambda)"

The distance from center to the tangent plane:

"d=\frac{|4\lambda+6\lambda-15|}{\sqrt{4^2+3^2}}=|2\lambda-3|"

We have:

"R=d"

Then:

"\sqrt{a+14\lambda^2}=|2\lambda-3|"

"a+14\lambda^2=4\lambda^2-12\lambda+9"

"10\lambda^2+12\lambda+a-9=0"

"\lambda=\frac{-12\pm\sqrt{144-40(a-9)}}{20}=\frac{-6\pm\sqrt{38-10(a-9)}}{10}"

Substitute "\lambda" in the equation of the sphere, and we get equations of two different spheres.